# Global Deep Forecasting with Patient-Specific Pharmacokinetics

### v8

@@ 3.3 Modeling Dose-Dependent Effects
A majority of drugs are assumed to adhere to linear PK, where it is assumed that there is a linear relationship between drug dose and concentration in the body (Rosenbaum, 2017). A characteristic of this assumption is that PK parameters are constant and do not change with dose. In practice, a drug’s PK parameters can vary depending on factors such as the drug’s properties, the type of injection, and the injection site. Nonlinear PK occurs when the relationship between drug dose and body concentration is not linear, leading to changes in PK parameters with dose. For instance, subcutaneous injections are administered into the subcutaneous tissue, where higher doses can saturate elimination pathways, causing slower absorption rates (Søeborg et al., 2009; Gradel et al., 2018). This variability in PK parameters can result in dose-dependent plasma drug concentration profiles.
Insulin pumps administer the same type of insulin (rapid-acting) for the basal and bolus doses. However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs. 12 and 13 in Appendix A.1. To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity, so that we can capture potential PK variations without resorting to a fully nonlinear PK model. This method simplifies the modeling process by maintaining linear assumptions within each dose category. It avoids the complexity while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics. Specifically, we separately estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values. For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
$\widecheck{\textbf{x}}^{(i)}_{\text{basal}}$ $=C(t,\textbf{x}^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)}),$ (6)
$\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C(t,\textbf{x}^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}).$ (7)
Treating each dose category as a distinct entity is an approximation and may not capture all nuances of dose-dependent kinetics. However, all physiological models approximate complex biological processes. Our goal is to develop models validated under established linear PK assumptions from the pharmacology literature while still accounting for dose-dependent effects, balancing simplicity and precision. This framework also ensures broader applicability across various medications, as most drugs are assumed to follow linear PK (Rosenbaum, 2017).
@@ 3.4 Models
Global models are trained on all patient data to create a single model. Global model baselines include: (1) univariate models trained with only the blood glucose signal and static features, (2) models trained with univariate blood glucose signals, static features, and sparse historical exogenous features (CHO and insulin dose values), and (3) models trained with the univariate blood glucose signal, static features, and sparse historical exogenous features transformed using the “Sum Total” (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022). Univariate models are included as baselines to highlight forecast improvements from training on historical exogenous variables, motivating advanced transformations for sparse, continuous exogenous features using methods like our PK encoder.
@@
Patient-specific sparse exogenous models as well as patient-specific exogenous models using the PK encoder were compared against our hybrid global-local PK encoder model.
@@ 3.5 Model Training
Models were trained using Adam optimizer (Kingma and Ba, 2017). Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected. Huber estimator was selected as the loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
@@
Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set. Optimal hyperparameters were selected based on a validation set with cross-validation. More information on model training and hyperparameters in provided in Appendix A.3.
@@ 3.6 Model Evaluation
Trained models were used to generate rolling window forecasts at one-step intervals. Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE):
We compute results across all points within the forecast horizon, and present the average results over 8 trials of individually trained models. We also evaluate the performance of the models exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation. To evaluate whether hybrid global-local PK models significantly outperform baselines, we use paired t-test statistics computed on the patients’ mean MAE results.
@@ 4 Cohort
We use simulated and real-world datasets containing patient insulin, CHO time series, and blood glucose measurements. These open-source datasets ensure reproducibility and are widely used in blood glucose forecasting research, featuring some of the largest patient cohorts (Xie and Wang, 2020; Rubin-Falcone et al., 2020, 2022; Rabby et al., 2021; Li et al., 2019b).
@@
Figure 5: The hybrid global-local NHITS-PK model trained on all OhioT1DM individuals (orange, solid median) has a lower average MAE than NHITS-PK local models (blue, dashed median).
@@
[…]
@@
w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
@@
Twelve de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a continuous glucose monitor (CGM) at a frequency of 5-minutes, basal rate information, bolus insulin (units), and CHO values. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. Missing values were forward filled to prevent data leakage. The sample count and percentages of missing data for each subject are included in Table 4 in Appendix A.1. Appendix A.1 also contains examples of blood glucose time series plots for 6 subjects, shown in Fig. 11, as well as basal and bolus dose distributions for the 12 patients, shown in Fig. 13.
@@ 5 Results
For simulated data, the TFT-PK model achieves the lowest MAE, on average, among all models for both evaluations across all models and critical values. The PK models, including TFT-PK, NBEATSx-PK, and NHITS-PK outperform their sparse exogenous model counterparts by up to 5.4%, 3.7%, 9.2%, respectively, on evaluations across all values and 7.7%, 6.6%, 16.4% on evaluations across critical values for individual subjects. For the OhioT1DM dataset, the NHITS-PK model achieves the lowest MAE, on average, among all models for both evaluations across all values and critical values. The PK models, including, TFT-PK, NBEATSx-PK, and NHITS-PK, outperform their sparse exogenous model counterparts by up to 5.2%, 8.5%, 4.9% on evaluations across critical values for individual patients. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Appendix C.3 Tables 8 and 9, respectively.
@@
For the simulated dataset, the TFT-PK, NBEATSx-PK, and NHITS-PK models achieve significantly lower forecast errors for subjects for evaluations on all values (p-values: 4.6e-4 [TFT-PK]; 7.9e-5 [NBEATSx-PK]; 2.0e-7 [NHITS-PK]) and critical values (p-values: 5.7e-4 [NBEATSx-PK]; 3.0e-5 [NHITS-PK]) compared to their sparse exogenous global model counterparts. For the OhioT1DM dataset, the NBEATSx-PK and NHITS-PK models also achieve significantly lower forecast errors for subjects for both evaluations on all values (p-values: 3.4e-3 [NBEATSx-PK]; 1.3e-3 [NHITS-PK]) and critical values (p-values: 4.4e-4 [NBEATSx-PK]; 7.0e-3 [NHITS-PK]) compared to their sparse exogenous global model counterparts. Additionally, Table 2 shows that TFT-PK and NHITS-PK models achieve significantly lower errors, on average—2.4% (p-value: 3.6e-12) and 2.3% (p-value: 4.1e-5), respectively—for subjects in evaluations across all values, compared to their “Sum Total” model counterparts.
@@
Table 2: PK models outperform Sum Total (ST) models for both Simulated and OhioT1DM datasets. Best result is in bold.
@@
Table 3: MAE computed across forecast horizon values for local NHITS, global NHITS, local NHITS-PK, and hybrid global-local NHITS-PK on the 12 subjects from the OhioT1DM dataset. Best results in bold.
@@
Figure 8: The NHITS-PK model captures insulin treatment effects: forecasts with removed bolus doses (yellow, dash-dotted) and augmented doses (red, dotted), result in higher and lower glucose predictions, respectively, compared to the original doses (orange, dashed). Each point represents the final forecasted timestamp from a rolling window forecast.
@@ 6 Discussion
The proposed PK encoder can substantially reduce forecasting errors in deep learning models for both simulated and real-world data. These improvements are not only statistically significant, but they have potential for positive real-world clinical impact. In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts these events earlier, up to 3.6 minutes compared to the NHITS sparse exogenous baseline model, for individual OhioT1DM patients, as shown in Fig. 9. This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations. Improving the prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009). Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing. While we demonstrate performance improvements across the top three models, our approach can be easily integrated into other deep learning architectures as new tools emerge. It can also be easily adapted for other types of medication administration, such as oral or intravenous injections.
@@
In addition to improving performance over global model counterparts, our hybrid global-local model architecture produces forecasts with significantly smaller errors than patient-specific models. This result is of substantial importance in demonstrating that global models can integrate cohort-level information with patient-specific parameters to produce more accurate personalized forecasts. Moreover, our proposed hybrid global-local model offers practical benefits beyond improved forecasting performance as it requires less computational resources and time to train than individual patient models. Training patient-specific NHITS-PK models took approximately seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space. The hybrid model also has the potential to be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
@@
Our study demonstrates that leveraging domain knowledge to integrate exogenous variables in models offers performance benefits over using sparse features or the “Sum Total” feature engineering approach, both of which are outperformed by our proposed deep PK encoding method. Furthermore, dose-dependent PK encoding strategies, such as those proposed in Section 3.3, can lead to more realistic models, as highlighted by our ablation study in Appendix C.2 where we demonstrate that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset.
@@
This study has two limitations related to the PK encoder: we assume that the concentration curves exhibit 100% bioavailability and maintain a consistent peak concentration across patients. These simplifying assumptions are intended to focus model learning on the parameter $k$, rather than on the area or peak of the concentration curve, to emphasize the spread or duration of the concentration curve and to reduce potential issues with parameter identification. Future work will explore expanding concentration curve parameters to estimate area and peak parameters to potentially enable more flexible concentration curves. We will also aim to apply our approach to additional cohorts of patients, such as those with T2DM, as more data become available, to further validate our findings. Future work will also explore modeling dose-dependent effects using more dynamic nonlinear PK frameworks.
@@
This work was partially supported by the NIH award R01HL141916. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.
Data and Code Availability

### v9

@@ 3.3 Modeling Dose-Dependent Effects
A majority of drugs are assumed to adhere to linear pharmacokinetics (PK), where a linear relationship is expected between drug dose and its concentration in the body (Rosenbaum, 2017). In practice, however, whether a drug’s PK parameters follow linear or nonlinear behavior can vary based on factors such as the drug’s physicochemical properties, the route of administration, and the injection site. A key characteristic of the linear PK assumption is that PK parameters remain constant and do not change with variables such as dose or time. Dose- or time-dependent nonlinearity occurs when PK parameters change nonproportionally with dose or time, respectively. In dose-dependent nonlinear PK, the relationship between drug dose and body concentration is no longer linear, resulting in dose-related changes in PK parameters. While insulin is not expected to exhibit classic time-dependent nonlinearity, it may demonstrate dose-dependent nonlinearity. For instance, subcutaneous injections are administered into the subcutaneous tissue, where higher doses can saturate elimination pathways, causing slower absorption rates (Søeborg et al., 2009; Gradel et al., 2018). When PK parameters are dose-dependent, the resulting plasma drug concentration profiles also vary with dose.
Insulin pumps administer the same type of insulin (rapid-acting) for the basal and bolus doses. However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs. 13 and 14 in Appendix A.1. To account for potential dose-dependent effects, we model basal and bolus doses separately by learning distinct PK parameters, $\mathbf{k}_{\text{basal}}$ and $\mathbf{k}_{\text{bolus}}$, allowing us to capture PK variations without resorting to a fully nonlinear model. This approach maintains linear assumptions within each dose type, simplifying the modeling process while still capturing differences between low-dose (basal) and high-dose (bolus) PK. For a single dose of basal insulin $\mathbf{x}_{\text{basal}}$ or bolus insulin $\mathbf{x}_{\text{bolus}}$, administered at time step $t_{d}$, we compute the plasma insulin concentration at time $t\geq t_{d}$ for patient $i$:
$\widecheck{\textbf{x}}^{(i)}_{\text{basal}}$ $=C\left(\nu(t-t_{d}),\textbf{x}^{(i)}_{t_{d},\text{basal}},\textbf{k}_{\text{basal}}^{(i)}\right),$ (6)
$\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C\left(\nu(t-t_{d}),\textbf{x}^{(i)}_{t_{d},\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}\right).$ (7)
Treating each dose category as a distinct entity is an approximation and may not capture all nuances of dose-dependent nonlinear PK. All physiological models are approximations of complex biology. Our goal is to evaluate a novel deep learning-based PK modeling approach based on established linear PK knowledge, while accounting for potential dose-dependent effects, balancing simplicity and precision. This framework also ensures broader applicability across various medications, as most drugs are assumed to follow linear PK (Rosenbaum, 2017).
@@ 3.4 Models
Global models are trained on all patient data to create a single model. Global model baselines include: (1) models trained with only the blood glucose signal and static features, (2) models trained with blood glucose signals, static features, and sparse historical exogenous features (CHO and insulin dose values), and (3) models trained with blood glucose signals, static features, and sparse historical exogenous features transformed using the “Sum Total” (ST) approach, which converts features into a cumulative sum of values across an input window (Rubin-Falcone et al., 2022). We include models trained without exogenous variables as baselines to show the impact of incorporating sparse exogenous information for each model. Together with models trained on sparse and Sum Total features, these baselines demonstrate the importance of effective transformations—such as our PK encoder—in reducing forecast error.
@@
Patient-specific models using sparse exogenous inputs, as well as those incorporating the PK encoder, were compared against our hybrid global-local PK encoder model. These baselines highlight the effectiveness of hybrid global-local architectures that leverage cross-learning from multiple time series.
@@ 3.5 Model Training
Models were trained using Adam optimizer (Kingma and Ba, 2017). Optimal hyperparameters were selected via cross-validation on a validation set by minimizing Huber loss across all prediction time points. Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set. Huber estimator was selected as the loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
@@
More information on model training and hyperparameters in provided in Appendix A.3.
@@ 3.6 Model Evaluation
Trained models were used to generate rolling window forecasts at one-step intervals. Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE). We compute results across all points within the forecast horizon, and present the average results over 8 trials of individually trained models. We also evaluate the performance of the models exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation. To evaluate whether hybrid global-local PK models significantly outperform baselines, we use paired t-tests computed on the patients’ mean MAE results. Metric formulations and statistical tests are detailed in Appendix A.4.
@@ 4 Cohort
We use simulated and real-world datasets containing patient insulin, CHO time series, and blood glucose measurements. These open-source datasets ensure reproducibility and are among the largest publicly available cohorts for blood glucose forecasting, making them widely used in current research within this domain (Xie and Wang, 2020; Rubin-Falcone et al., 2020, 2022; Rabby et al., 2021; Li et al., 2019b).
@@
Figure 5: The hybrid global-local NHITS-PK model trained on all OhioT1DM individuals (orange, solid median) has a lower MAE than NHITS-PK local models (blue, dashed median).
@@
Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose $\leq 70~|\geq 180$). Models are separated with horizontal lines into statistical, linear, RNN-based, CNN-based, Transformer-based, and MLP-based, and PK model groups, respectively. Result for models with (w/) and without (w/o) exogenous (exog.) variables are shown where applicable. Best results are highlighted in bold, second best results are underlined. PK models with improved forecasts over their respective exogenous baseline model are highlighted in blue.
@@
w/o Exog.
w/ Exog.
w/o Exog.
w/ Exog.
w/o Exog.
w/ Exog.
w/o Exog.
w/ Exog.
@@
Twelve de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a continuous glucose monitor (CGM) at a frequency of 5-minutes, basal rate information, bolus insulin (units), and CHO values. The maximum number of test samples consistent across all patients (2691 timestamps) was used to obtain equal test sets of approximately 9 days. Missing values were forward filled to prevent data leakage. Sample counts and missing data percentages for each subject are included in Table 4 in Appendix A.1. Appendix A.1 also contains examples of blood glucose time series plots for 6 subjects, shown in Fig. 12, as well as basal and bolus dose distributions for the 12 patients, shown in Fig. 14.
@@ 5 Results
For simulated data, the TFT-PK model achieves the lowest MAE, on average, among all models for both evaluations across all models and critical values. The PK models, including TFT-PK, NBEATSx-PK, and NHITS-PK outperform their sparse exogenous model counterparts by up to 5.4%, 3.7%, 9.2%, respectively, on evaluations across all values and 7.7%, 6.6%, 16.4% on evaluations across critical values for individual patients. For the OhioT1DM dataset, the NHITS-PK model achieves the lowest MAE, on average, among all models for both evaluations across all values and critical values. The PK models, including, TFT-PK, NBEATSx-PK, and NHITS-PK, outperform their sparse exogenous model counterparts by up to 5.2%, 8.5%, 4.9% on evaluations across critical values for individual patients. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Appendix C.7 Tables 11 and 12.
@@
For the simulated dataset, the TFT-PK, NBEATSx-PK, and NHITS-PK models achieve significantly lower forecast errors for patients for evaluations on all values (p-values: 4.6e-4 [TFT-PK]; 7.9e-5 [NBEATSx-PK]; 2.0e-7 [NHITS-PK]) and critical values (p-values: 5.7e-4 [NBEATSx-PK]; 3.0e-5 [NHITS-PK]) compared to their sparse exogenous global model counterparts. For the OhioT1DM dataset, the NBEATSx-PK and NHITS-PK models also achieve significantly lower forecast errors for patients for both evaluations on all values (p-values: 3.4e-3 [NBEATSx-PK]; 1.3e-3 [NHITS-PK]) and critical values (p-values: 4.4e-4 [NBEATSx-PK]; 7.0e-3 [NHITS-PK]) compared to their sparse exogenous global model counterparts. The TFT-PK model does not outperform its sparse exogenous counterpart for L=120L\!=\!120. However, for L=180L\!=\!180, it achieves significantly lower error (p-value: 4.8e-2) than its sparse exogenous variant and records the lowest error among all TFT model variants, as shown in Table 8. Additionally, Table 2 shows that TFT-PK and NHITS-PK models achieve significantly lower errors, on average—2.4% (p-value: 3.6e-12) and 2.3% (p-value: 4.1e-5), respectively—for subjects in evaluations across all values, compared to their “Sum Total” model counterparts.
@@
Table 2: PK models outperform Sum Total (ST) models for both Simulated and OhioT1DM datasets. The best results are in bold.
@@
Table 3: MAE computed across forecast horizon values for local NHITS, global NHITS, local NHITS-PK, and hybrid global-local NHITS-PK on the 12 subjects from the OhioT1DM dataset. The best results are in bold.
@@ 6 Discussion
The proposed PK encoder can substantially reduce forecasting errors in deep learning models for both simulated and real-world data. In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts these events earlier, up to 3.6 minutes compared to the NHITS sparse exogenous baseline model, for individual OhioT1DM patients, as shown in Fig. 10. The TFT-PK model achieves a true positive rate of 0.96 on the simulated dataset, while the NHITS-PK model reaches 0.89 on the OhioT1DM dataset, showcasing their effectiveness in predicting critical events while leaving room for further improvement. A complete analysis of hyper- and hypoglycemic event prediction is provided in Appendix C.4.
Figure 8: The NHITS-PK model captures insulin treatment effects: forecasts with removed bolus doses (yellow, dash-dotted) and augmented doses (red, dotted), result in higher and lower glucose predictions, respectively, compared to the original doses (orange, dashed). Each point represents the final forecasted time step from a rolling window forecast.
@@
In addition to improving performance over global model baselines, our hybrid global-local model architecture produces forecasts with significantly smaller errors than patient-specific models. This result is important in demonstrating that global models can integrate cohort-level information with patient-specific parameters to produce more accurate personalized forecasts. Evaluations on the OhioT1DM dataset also demonstrate that our hybrid global-local model remains effective even with 20% of missing timestamps in the training set for individual patients, as shown in Table 3, with missing data detailed in Table 4. Moreover, our proposed hybrid global-local model offers practical benefits beyond improved forecasting performance as it requires less computational resources and time to train than individual patient models. Training patient-specific NHITS-PK models took approximately seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space. Additionally, the PK models are as computationally efficient as or more efficient than their sparse exogenous counterparts in terms of floating-point operations per second (FLOPs) and the number of trainable parameters, as shown in Table 9 in Appendix C.5.
(a) Simulated Dataset
(b) OhioT1DM Dataset
(c) Simulated Dataset
(d) OhioT1DM Dataset
Figure 9: True positive rate (TPR) for the Simulated and OhioT1DM datasets at various glucose thresholds for (a, b) hyperglycemia (blood glucose $\geq 180$) and (c, d) hypoglyemia (blood glucose $\leq 70$) event warnings. TPR represents the proportion of forecast horizon windows in which the model correctly predicts an event when one actually occurs. The best performing TFT-PK model for the Simulated dataset predicts hyperglycemic and hypoglycemic events in forecast windows at an approximate TPR of 0.96 and 0.93, respectively. The best performing NHITS-PK model for the OhioT1DM dataset predicts hyperglycemic and hypoglycemic events in forecast windows at an approximate TPR of 0.89 and 0.69, respectively.
@@
Our study demonstrates that leveraging PK domain knowledge to integrate exogenous variables in models offers performance improvements over using sparse features or the “Sum Total” feature engineering approach. Furthermore, dose-dependent PK encoding strategies, such as those proposed in Section 3.3, can lead to more realistic models, as highlighted by our ablation study in Appendix C.2 where we demonstrate that using our PK encoder to infer the same k value for both bolus and basal insulin, rather than separate values, results in worse performance on the OhioT1DM dataset.
Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing. We discuss the feasibility of integrating such deep learning models into clinical decision-support systems in section C.6. While we demonstrate performance improvements across the top three models, our approach can be easily integrated into other deep learning architectures as new tools emerge. It can also be adapted for other types of medication administration, such as oral or intravenous injections.
@@
This study has two limitations related to the PK encoder: we assume 100% bioavailability and a consistent peak concentration across patients. These assumptions focus model learning on the parameter $k$, rather than the area or peak of the concentration curve, to help mitigate potential issues with parameter identification. Future work will explore expanding concentration curve parameters as well as using nonlinear PK frameworks to potentially enable more flexible concentration curves. The proposed method is demonstrated on T1DM patient data but has potential to be adapted to other forecasting tasks and populations, such as insulin-dependent T2DM patients, with appropriate cohort data, PK model adjustments, and model training. Future work will apply our approach to additional cohorts, including T2DM, as more data becomes available to further validate our findings. Future work will also explore methods for estimating patient-specific parameters in a global model for new patients with limited or no historical data.
@@
This work has been partially supported by the NIH awards R01NR013912 and R01DK131586, and NSF awards 2406231 and 2427948. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.